ok i am in algerbra and my final is coming up in a couple of weeks and i honestly have no clue how to do mixtrue problems, is there any way you can help????
Found 2 solutions by edjones, josmiceli:
Answer by edjones(8007)

Try purplemath.com

Answer by josmiceli(19441)

Here's a sample problem that I solved:
-------------------------------
Soybean is 18% protein
Cornmeal is 9% protein
How many lbs of each should be mixed together to get 360-lb mixture that is 12% protein?
-----------------------------
Let a = pounds of soybean needed
Let b = pounds of cornmeal needed
given:
.18a = protein in soybeans
.09b = protein in cornmeal
(1) a + b = 360
(2) ( .18a + .09b ) / 360 = .12
-----------------------
(2) .18a + .09b = .12*360
(2) .18a + .09b = 43.2
(2) 18a + 9b = 4320
Multiply both sides of (1) by 9 and
subtract (1) from (2)
(2) 18a + 9b = 4320
(1) -9a - 9b = -3240
9a = 1080
a = 120
and
(1) a + b = 360
b = 360 - 120
b = 240
120 pounds of soybeans are needed
240 pounds of cornmeal are needed
check answer:
(2) ( .18a + .09b ) / 360 = .12
(2) .18*120 + .09*240 = .12*360
(2) 21.6 + 21.6 = 43.2
(2) 43.2 = 43.2
OK